OpenStudy (anonymous):

What is a number between 80 and 100 with at least 3 factors and one of them is 5? HELP

OpenStudy (anonymous):

someone help me

OpenStudy (acxbox22):

first list all the numbers between 80 and 100 that can be divided by 5

OpenStudy (anonymous):

85 90 95

OpenStudy (phi):

divide each by 5 (to simplify them) what do you get ?

OpenStudy (acxbox22):

^^ do that to get the other factors of 85,90, and 95

OpenStudy (anonymous):

85/5=17 90/5=18 95/5=19

OpenStudy (phi):

yes, that is my idea

OpenStudy (phi):

now look for a number that has at least 2 factors notice two of them are prime (no more factors, except themselves and 1 )

OpenStudy (anonymous):

im confused

OpenStudy (phi):

do you agree 17 is prime ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

Here is a list of prime numbers http://metricconversion.biz/list-of-first-100-prime-numbers.html primes do not factor

OpenStudy (anonymous):

and so is 19

OpenStudy (phi):

notice 85= 5*17 both 5 and 17 are prime numbers. that means they cannot be factored. in other words, 85 has exactly two factors , namely 5 and 17 but you want a number with at least 3 factors. so cross 85 off your list as a possible answer.

OpenStudy (phi):

yes, which leaves 90= 5*18 but 18 is not prime, so it can be factored